Sam is building a cabinet. The materials that he needs cost $480. To determine the price of the cabinet for his customer, Sam charges 30% more than the cost of the materials. Sam asks for a downpayment of 40% of the total before the work begins. The costumer pays the balance when the job is finished.
Answer by nerdybill(7384)

Sam is building a cabinet. The materials that he needs cost $480. To determine the price of the cabinet for his customer, Sam charges 30% more than the cost of the materials. Sam asks for a downpayment of 40% of the total before the work begins. The costumer pays the balance when the job is finished.
.
Price of the cabinet (to the customer) is:
"cost of materials" + "30% of cost of materials"
480 + .30(480)
= 480(1+.30)
= 480(1.30)
= $624 (cost of cabinet)
.
Down payment (due at start of work):
.40(624) = $249.60
.
Amount due at completion:
624-249.60 = $374.4